OpenStudy (anonymous):

The slope m and the y-intercept b of the line 10x + y + 8 = 0 are:

OpenStudy (jdoe0001):

solve " 10x + y + 8 = 0 " for "y" see what you get

OpenStudy (anonymous):

-10x-8 ?

OpenStudy (jadeishere):

y = -10x - 8 So -10 = slope and -8 = yintercept y = mx + b

jigglypuff314 (jigglypuff314):

yes! :) then y = mx + b where m = slope and b = y-intercept y = -10x+(-8) y = m x + b

OpenStudy (anonymous):

YEach so your slope is -10 and intercept = -8

OpenStudy (anonymous):

Yes*